An ampoule contains a solution of drug of 150micrograms per 2.5mililitres.convert this dose to grams per litre

1
Expert's answer
2022-04-11T14:00:49-0400
"\frac{150\:\rm \mu g}{2.5\:\rm mL}=\frac{150*10^{-6}\:\rm g}{2.5*10^{-3}\:\rm L}=0.06\: \rm g/L"